# Bounding the SNPR distance between two tree-child networks using generalised agreement forests

### v1

@@ 6. Bounding the tree-child SNPR distance
Next suppose that there exists an element $i\in\{0,1,2,\ldots,t-2\}$ such that ${\mathcal{N}}_{i+1}$ is obtained from ${\mathcal{N}}_{i}$ by an SNPR± and ${\mathcal{N}}_{i+2}$ is obtained from ${\mathcal{N}}_{i+1}$ by an SNPR-. Let $e=(u,v)$ be the edge of ${\mathcal{N}}_{i}$ that is deleted in the process of obtaining ${\mathcal{N}}_{i+1}$. Furthermore, let $p_{u}$ be the parent of $u$ and let $c_{u}$ be the child of $u$ that is not $v$ in ${\mathcal{N}}_{i}$. Since $u$ is a tree vertex, $p_{u}$ and $c_{u}$ are well defined. Let $f=(p_{u^{\prime}},c_{u^{\prime}})$ be the edge of the digraph resulting from ${\mathcal{N}}_{i}$ by deleting $e$ and suppressing $u$ that is subdivided with a new vertex $u^{\prime}$ such that $(u^{\prime},v)$ is an edge in ${\mathcal{N}}_{i+1}$. If $p_{u^{\prime}}=p_{u}$ and $c_{u^{\prime}}=c_{u}$, then ${\mathcal{N}}_{i}\cong{\mathcal{N}}_{i+1}$, and so (iii) holds. Hence, we may assume that $p_{u^{\prime}}\neq p_{u}$ or $c_{u^{\prime}}\neq c_{u}$, and so $(p_{u},c_{u})$ is an edge in ${\mathcal{N}}_{i+1}$. Let $e^{\prime}$ be the edge that is deleted in obtaining ${\mathcal{N}}_{i+2}$ from ${\mathcal{N}}_{i+1}$. There are three cases to consider. First assume that $e^{\prime}=(p_{u},c_{u})$. Then $(u,c_{u})$ is a reticulation edge in ${\mathcal{N}}_{i}$. Let $w$ be the parent of $c_{u}$ that is not $u$ in ${\mathcal{N}}_{i}$. Obtain ${\mathcal{N}}_{i+1}^{\prime}$ by deleting $(u,c_{u})$ and suppressing the two resulting degree-two vertices. By Lemma 2.2, ${\mathcal{N}}_{i+1}^{\prime}$ is tree-child. Moreover, noting that $(p_{u},v)$ is an edge in ${\mathcal{N}}_{i+1}^{\prime}$, it follows that ${\mathcal{N}}_{i+2}$ can be obtained from ${\mathcal{N}}_{i+1}^{\prime}$ by an SNPR± that deletes $(p_{u},v)$, suppresses $p_{u}$, subdivides $(p_{u^{\prime}},c_{u^{\prime}})$ if $c_{u}\neq p_{u^{\prime}}$ (resp. subdivides $(w,c_{u^{\prime}})$ if $c_{u}=p_{u^{\prime}}$) with a new vertex $u^{\prime}$, and adds the edge $(u^{\prime},v)$, and so (iv) holds. Second assume that $e^{\prime}=(u^{\prime},c_{u^{\prime}})$. Then $f$ is a reticulation edge in ${\mathcal{N}}_{i}$. Noting that $p_{u^{\prime}}$ is a tree vertex in ${\mathcal{N}}_{i}$, let $s$ be the parent of $p_{u^{\prime}}$, and let $t$ be the child of $p_{u^{\prime}}$ that is not $c_{u^{\prime}}$. Now obtain ${\mathcal{N}}_{i+1}^{\prime}$ from ${\mathcal{N}}_{i}$ by deleting $f$ and suppressing the two resulting degree-two vertices. Again by Lemma 2.2, ${\mathcal{N}}_{i+1}^{\prime}$ is tree-child. Furthermore, $(s,t)$ is an edge in ${\mathcal{N}}_{i+1}^{\prime}$. Then obtain ${\mathcal{N}}_{i+2}$ from ${\mathcal{N}}_{i+1}^{\prime}$ by an SNPR± that deletes $(u,v)$, suppresses $u$, subdivides $(s,t)$ with a new edge $u^{\prime}$, and adds the edge $(u^{\prime},v)$. Again (iv) holds. Third assume that $e^{\prime}\neq(p_{u},c_{u})$ and $e^{\prime}\neq(u^{\prime},c_{u^{\prime}})$. Then $e^{\prime}$ is an edge of ${\mathcal{N}}_{i}$. Let ${\mathcal{N}}^{\prime}_{i+1}$ be the tree-child network obtained from ${\mathcal{N}}_{i}$ by deleting $e^{\prime}$ and suppressing the two resulting degree-two vertices. Since $e$ and $f$ are edges of ${\mathcal{N}}_{i+1}^{\prime}$, it now follows that ${\mathcal{N}}_{i+2}$ can be obtained from ${\mathcal{N}}_{i+1}^{\prime}$ by an SNPR± that deletes $e$, suppresses $u$, subdivides $f$ with a new vertex $u^{\prime}$, and adds the edge $(u^{\prime},v)$. Again, (iv) holds, thereby completing the proof of the lemma. ∎
Corollary 6.3.
@@
The corollary follows from repeated applications of Lemma 6.2. ∎
The proof of Theorem 1.1 is an amalgamation of the next two lemmas.
Lemma 6.4.
@@
Let ${\mathcal{D}}=\{D_{\rho},D_{1},D_{2},\ldots,D_{k}\}$ be an agreement tree-child digraph for ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$, and let =${\mathcal{R}}=\{R_{\rho},R_{1},R_{2},\ldots,R_{k}\}$ and =${\mathcal{R}}^{\prime}=\{R_{\rho}^{\prime},R_{1}^{\prime},R_{2}^{\prime},\dots,R_{k}^{\prime}\}$ be an extension of ${\mathcal{D}}$ in ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$, respectively. Furthermore, let $c_{\mathcal{D}}=|E_{\mathcal{N}}-E_{\mathcal{R}}|$ and $c_{\mathcal{D}}^{\prime}=|E_{{\mathcal{N}}^{\prime}}-E_{{\mathcal{R}}^{\prime}}|$ be the cut size of ${\mathcal{D}}$ in ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$, respectively. By Lemma 4.4, we may assume that ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ is a root extension of ${\mathcal{D}}$ in ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$, respectively. We show by induction on $c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}$ that $d_{{\rm tc}}({\mathcal{N}},{\mathcal{N}}’)\leq c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}$. The lemma then follows by choosing ${\mathcal{D}}$ to be an agreement tree-child digraph for ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ such that $c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}=m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$.
@@
If $c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}=0$, then ${\mathcal{N}}\cong{\mathcal{N}}^{\prime}$ and consequently $d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})=0$. Hence, the result follows. Now assume that $c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}\geq 1$ and that the result holds for all pairs of tree-child networks ${\mathcal{N}}_{1}$ and ${\mathcal{N}}_{1}^{\prime}$ on the same leaf set that have an agreement tree-child digraph ${\mathcal{D}}_{1}$ with cut size $c_{{\mathcal{D}}_{1}}$ and $c_{{\mathcal{D}}_{1}}^{\prime}$ of ${\mathcal{D}}_{1}$ in ${\mathcal{N}}_{1}$ and ${\mathcal{N}}_{1}^{\prime}$, respectively, such that $c_{{\mathcal{D}}_{1}}+c_{{\mathcal{D}}_{1}}^{\prime}<c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}$.
@@
We first establish the lemma for a case that is easy to deal with. To this end, assume that there exists a reticulation edge $e=(u,v)$ in ${\mathcal{N}}$ or ${\mathcal{N}}^{\prime}$ that is not contained in ${\mathcal{R}}$ or ${\mathcal{R}}^{\prime}$, respectively. Without loss of generality, we may assume that $e\in{\mathcal{N}}$. As ${\mathcal{N}}$ is tree-child, $u$ is a tree vertex. Let ${\mathcal{N}}^{\prime\prime}$ be the phylogenetic network obtained from ${\mathcal{N}}$ by an SNPR- that deletes $e$ and suppresses the two resulting degree-two vertices. By Lemma 2.2, ${\mathcal{N}}^{\prime\prime}$ is tree-child. Furthermore, by Lemma 6.1(i), ${\mathcal{D}}$ is a tree-child digraph for ${\mathcal{N}}^{\prime\prime}$ and there exists a root extension ${\mathcal{R}}^{\prime\prime}$ of ${\mathcal{D}}$ in ${\mathcal{N}}^{\prime\prime}$ such that $c_{\mathcal{D}}-1=c_{\mathcal{D}}^{\prime\prime}$, where $c_{\mathcal{D}}^{\prime\prime}=|E_{{\mathcal{N}}^{\prime\prime}}-E_{{\mathcal{R}}^{\prime\prime}}|.$ Since $c_{\mathcal{D}}^{\prime\prime}+c_{\mathcal{D}}^{\prime}<c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}$, it now follows from the induction assumption that
$d_{\rm tc}({\mathcal{N}}^{\prime\prime},{\mathcal{N}}^{\prime})\leq c_{\mathcal{D}}^{\prime\prime}+c_{\mathcal{D}}^{\prime}.$
@@
Hence, there exists a tree-child SNPR sequence $\sigma$ connecting ${\mathcal{N}}^{\prime\prime}$ and ${\mathcal{N}}^{\prime}$ with $w(\sigma)\leq c_{\mathcal{D}}^{\prime\prime}+c_{\mathcal{D}}^{\prime}$. Moreover, since ${\mathcal{N}}^{\prime\prime}$ can be obtained from ${\mathcal{N}}$ by a single SNPR-, we have
$d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})\leq 1+c_{\mathcal{D}}^{\prime\prime}+c_{\mathcal{D}}^{\prime}=c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime},$
@@
Hence, by symmetry, there exists a tree edge $e=(u,w)$ in ${\mathcal{N}}$ that is not in ${\mathcal{R}}$. Choose $e$ such that each directed path from $w$ to a leaf in ${\mathcal{N}}$ only consists of edges in ${\mathcal{R}}$. Since ${\mathcal{N}}$ is acyclic such an edge exists. Let $D_{i}$ be the element in ${\mathcal{D}}$ with $i\in\{\rho,1,2,\ldots,k\}$ such that the root extension $R_{i}$ of $D_{i}$ in ${\mathcal{N}}$ contains $w$. By the choice of $e$, $R_{i}$ exists and each edge in ${\mathcal{N}}$ that lies on a directed path from $w$ to a leaf is an element of $R_{i}$. Thus, if $u=\rho$, then ${\mathcal{N}}\cong{\mathcal{N}}^{\prime}$ and the result follows as $d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})=0$. We may therefore assume that $u\neq\rho$. The next statement follows from Lemma 4.1(i), the additional assumption that $u\neq\rho$, and assumption (A).
6.4.1.
@@
By the choice of $e$, observe that the root path of $w$ consists only of $w$. In turn, $w$ has in-degree zero and out-degree zero or two in $R_{i}$. Hence, $w$ corresponds to a unique vertex, say $w_{\mathcal{D}}$, of $D_{i}$. Let $r_{w^{\prime}}$ be the vertex in ${\mathcal{N}}^{\prime}$ that $w_{\mathcal{D}}$ corresponds to. Now consider the root extension $R_{i}^{\prime}$ of $D_{i}$ in ${\mathcal{N}}^{\prime}$. Let $w^{\prime}$ be the first vertex of the root path of $r_{w^{\prime}}$. By (A), we have that $w^{\prime}$ is a leaf, or has in-degree one and out-degree two in ${\mathcal{N}}^{\prime}$. Let $v^{\prime}$ be the parent of $w^{\prime}$ in ${\mathcal{N}}^{\prime}$. The next statement follows from Lemma 4.1(i).
6.4.2.
@@
Let $D_{j}$ be the element in ${\mathcal{D}}$ with $j\in\{\rho,1,2,\ldots,k\}$ such that the root extension $R_{j}^{\prime}$ of $D_{j}$ in ${\mathcal{N}}^{\prime}$ contains $v^{\prime}$. By Lemma 4.1(ii), $R_{j}^{\prime}$ exists. We may have $i=j$. We next construct a digraph ${\mathcal{D}}^{\prime}$ from ${\mathcal{D}}$ and a network ${\mathcal{N}}^{\prime\prime}$ from ${\mathcal{N}}$. After detailing the construction, we show that ${\mathcal{N}}^{\prime\prime}$ is a tree-child network that can be obtained from ${\mathcal{N}}$ by a single SNPR±, and that ${\mathcal{D}}^{\prime}$ is an agreement tree-child digraph for ${\mathcal{N}}^{\prime\prime}$ and ${\mathcal{N}}^{\prime}$. Guided by the second part of (6.4.2) and noting that, if $v^{\prime}=\rho$, then $\rho$ is a singleton component in ${\mathcal{D}}$, there are three cases to consider:
@@
({$({\mathcal{R}}^{\prime}-\{R_{i}^{\prime},R_{j}^{\prime}\})\cup\{R_{ij}^{\prime}\}$
@@
We next turn to ${\mathcal{D}}^{\prime}$ and show that ${\mathcal{D}}^{\prime}$ is tree-child. Since ${\mathcal{D}}$ is tree-child, it follows from the definition that ${\mathcal{R}}^{\prime}$ is tree-child. Moreover, since $w^{\prime}$ has in-degree zero in $R_{i}^{\prime}$, it follows that $w^{\prime}$ has in-degree one in $R_{ij}^{\prime}$. It is now straightforward to check that ({$({\mathcal{R}}^{\prime}-\{R_{i}^{\prime},R_{j}^{\prime}\})\cup\{R_{ij}^{\prime}\}$ is tree-child. Hence, again by definition, ${\mathcal{D}}^{\prime}$ is also tree-child.
@@
It now follows from the construction of ${\mathcal{N}}^{\prime\prime}$ from ${\mathcal{N}}$ and (6.4.4) that ${\mathcal{N}}^{\prime\prime}$ is a phylogenetic network on $X$. For the remainder of the proof, let $D_{u}$ be the element in ${\mathcal{D}}$ with $u\in\{\rho,1,2,\ldots,k\}$ such that the root extension $R_{u}$ of $D_{u}$ in ${\mathcal{N}}$ contains $u$. By Lemma 4.1(ii), $R_{u}$ exists.
6.4.5.
@@
Again using the same notation as in the construction of ${\mathcal{N}}^{\prime\prime}$ from ${\mathcal{N}}$, it follows from (6.4.1) that the newly added edge $(p,w)$ in ${\mathcal{N}}^{\prime\prime}$ is a tree edge. Noting that $u$ is a tree vertex by (6.4.1) in ${\mathcal{N}}$, let $p_{u}$ be the parent of $u$, and let $c_{u}$ be the child of $u$ that is not $w$ in ${\mathcal{N}}$. Observe that $(p_{u},c_{u})$ is an edge in ${\mathcal{N}}^{\prime\prime}$. Now assume that ${\mathcal{N}}^{\prime\prime}$ is not tree-child.

### v2

@@ 6 Bounding the tree-child SNPR distance
Next suppose that there exists an element $i\in\{0,1,2,\ldots,t-2\}$ such that ${\mathcal{N}}_{i+1}$ is obtained from ${\mathcal{N}}_{i}$ by an SNPR± and ${\mathcal{N}}_{i+2}$ is obtained from ${\mathcal{N}}_{i+1}$ by an SNPR-. Let $e=(u,v)$ be the edge of ${\mathcal{N}}_{i}$ that is deleted in the process of obtaining ${\mathcal{N}}_{i+1}$. Furthermore, let $p_{u}$ be the parent of $u$ and let $c_{u}$ be the child of $u$ that is not $v$ in ${\mathcal{N}}_{i}$. Since $u$ is a tree vertex, $p_{u}$ and $c_{u}$ are well defined. Let $f=(p_{u^{\prime}},c_{u^{\prime}})$ be the edge of the digraph resulting from ${\mathcal{N}}_{i}$ by deleting $e$ and suppressing $u$ that is subdivided with a new vertex $u^{\prime}$ such that $(u^{\prime},v)$ is an edge in ${\mathcal{N}}_{i+1}$. If $p_{u^{\prime}}=p_{u}$ and $c_{u^{\prime}}=c_{u}$, then ${\mathcal{N}}_{i}\cong{\mathcal{N}}_{i+1}$, and so (iii) holds. Hence, we may assume that $p_{u^{\prime}}\neq p_{u}$ or $c_{u^{\prime}}\neq c_{u}$, and so $(p_{u},c_{u})$ is an edge in ${\mathcal{N}}_{i+1}$. Let $e^{\prime}$ be the edge that is deleted in obtaining ${\mathcal{N}}_{i+2}$ from ${\mathcal{N}}_{i+1}$. There are three cases to consider, which are illustrated in Figure 7. First assume that $e^{\prime}=(p_{u},c_{u})$. Then $(u,c_{u})$ is a reticulation edge in ${\mathcal{N}}_{i}$. Let $w$ be the parent of $c_{u}$ that is not $u$ in ${\mathcal{N}}_{i}$. Obtain ${\mathcal{N}}_{i+1}^{\prime}$ by deleting $(u,c_{u})$ and suppressing the two resulting degree-two vertices. By Lemma 4, ${\mathcal{N}}_{i+1}^{\prime}$ is tree-child. Moreover, noting that $(p_{u},v)$ is an edge in ${\mathcal{N}}_{i+1}^{\prime}$, it follows that ${\mathcal{N}}_{i+2}$ can be obtained from ${\mathcal{N}}_{i+1}^{\prime}$ by an SNPR± that deletes $(p_{u},v)$, suppresses $p_{u}$, subdivides $(p_{u^{\prime}},c_{u^{\prime}})$ if $c_{u}\neq p_{u^{\prime}}$ (resp. subdivides $(w,c_{u^{\prime}})$ if $c_{u}=p_{u^{\prime}}$) with a new vertex $u^{\prime}$, and adds the edge $(u^{\prime},v)$, and so (iv) holds. Second assume that $e^{\prime}=(u^{\prime},c_{u^{\prime}})$. Then $f$ is a reticulation edge in ${\mathcal{N}}_{i}$. Noting that $p_{u^{\prime}}$ is a tree vertex in ${\mathcal{N}}_{i}$, let $s$ be the parent of $p_{u^{\prime}}$, and let $t$ be the child of $p_{u^{\prime}}$ that is not $c_{u^{\prime}}$. Now obtain ${\mathcal{N}}_{i+1}^{\prime}$ from ${\mathcal{N}}_{i}$ by deleting $f$ and suppressing the two resulting degree-two vertices, one of which is $p_{u^{\prime}}$. Again by Lemma 4, ${\mathcal{N}}_{i+1}^{\prime}$ is tree-child. Furthermore, $(s,t)$ is an edge in ${\mathcal{N}}_{i+1}^{\prime}$. Then obtain ${\mathcal{N}}_{i+2}$ from ${\mathcal{N}}_{i+1}^{\prime}$ by an SNPR± that deletes $(u,v)$, suppresses $u$, subdivides $(s,t)$ with a new vertex $p_{u^{\prime}}$, and adds the edge $(p_{u^{\prime}},v)$. Again (iv) holds. Third assume that $e^{\prime}\neq(p_{u},c_{u})$ and $e^{\prime}\neq(u^{\prime},c_{u^{\prime}})$. Then $e^{\prime}$ is an edge of ${\mathcal{N}}_{i}$. Let ${\mathcal{N}}^{\prime}_{i+1}$ be the tree-child network obtained from ${\mathcal{N}}_{i}$ by deleting $e^{\prime}$ and suppressing the two resulting degree-two vertices. Since $e$ and $f$ are edges of ${\mathcal{N}}_{i+1}^{\prime}$, it now follows that ${\mathcal{N}}_{i+2}$ can be obtained from ${\mathcal{N}}_{i+1}^{\prime}$ by an SNPR± that deletes $e$, suppresses $u$, subdivides $f$ with a new vertex $u^{\prime}$, and adds the edge $(u^{\prime},v)$. Again, (iv) holds, thereby completing the proof of the lemma. ∎
Corollary 12.
@@
The corollary follows from repeated applications of Lemma 11. ∎
The proof of Theorem 1 is an amalgamation of the next two lemmas.
Lemma 13.
@@
Let ${\mathcal{D}}=\{D_{\rho},D_{1},D_{2},\ldots,D_{k}\}$ be an agreement tree-child digraph for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$, and let ${\mathcal{R}}=\{R_{\rho},R_{1},R_{2},\ldots,R_{k}\}$ and ${\mathcal{R}}^{\prime}=\{R_{\rho}^{\prime},R_{1}^{\prime},R_{2}^{\prime},\dots,R_{k}^{\prime}\}$ be an extension of $\mathcal{D}$} in $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$, respectively. Furthermore, let $c_{\mathcal{D}}=|E_{\mathcal{N}}-E_{\mathcal{R}}|$ and $c_{\mathcal{D}}^{\prime}=|E_{{\mathcal{N}}^{\prime}}-E_{{\mathcal{R}}^{\prime}}|$ be the cut size of $\mathcal{D}$} in $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$, respectively. By Lemma 8, we may assume that $\mathcal{R}$} and $\mathcal{R}$$^{\prime}$ is a root extension of $\mathcal{D}$} in $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$, respectively. We show by induction on $c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}$ that $d_{{\rm tc}}({\mathcal{N}},{\mathcal{N}}’)\leqslant c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}$. The lemma then follows by choosing $\mathcal{D}$} to be an agreement tree-child digraph for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ such that $c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}=m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$.
@@
If $c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}=0$, then ${\mathcal{N}}\cong{\mathcal{N}}^{\prime}$ and consequently $d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})=0$. Hence, the result follows. Now assume that $c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}\geqslant 1$ and that the result holds for all pairs of tree-child networks ${\mathcal{N}}_{1}$ and ${\mathcal{N}}_{1}^{\prime}$ on the same leaf set that have an agreement tree-child digraph ${\mathcal{D}}_{1}$ with cut size $c_{{\mathcal{D}}_{1}}$ and $c_{{\mathcal{D}}_{1}}^{\prime}$ of ${\mathcal{D}}_{1}$ in ${\mathcal{N}}_{1}$ and ${\mathcal{N}}_{1}^{\prime}$, respectively, such that $c_{{\mathcal{D}}_{1}}+c_{{\mathcal{D}}_{1}}^{\prime}<c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}$.
@@
We first establish the lemma for a case that is easy to deal with. To this end, assume that there exists a reticulation edge $e=(u,v)$ in $\mathcal{N}$} or $\mathcal{N}$$^{\prime}$ that is not contained in $\mathcal{R}$} or $\mathcal{R}$$^{\prime}$, respectively. Without loss of generality, we may assume that $e\in{\mathcal{N}}$. As $\mathcal{N}$} is tree-child, $u$ is a tree vertex. Let $\mathcal{N}$$^{\prime\prime}$ be the phylogenetic network obtained from $\mathcal{N}$} by an SNPR- that deletes $e$ and suppresses the two resulting degree-two vertices. By Lemma 4, $\mathcal{N}$$^{\prime\prime}$ is tree-child. Furthermore, by Lemma 10(i), $\mathcal{D}$} is a tree-child digraph for $\mathcal{N}$$^{\prime\prime}$ and there exists a root extension $\mathcal{R}$$^{\prime\prime}$ of $\mathcal{D}$} in $\mathcal{N}$$^{\prime\prime}$ such that $c_{\mathcal{D}}-1=c_{\mathcal{D}}^{\prime\prime}$, where $c_{\mathcal{D}}^{\prime\prime}=|E_{{\mathcal{N}}^{\prime\prime}}-E_{{\mathcal{R}}^{\prime\prime}}|.$ Since $c_{\mathcal{D}}^{\prime\prime}+c_{\mathcal{D}}^{\prime}<c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}$, it now follows from the induction assumption that
$d_{\rm tc}({\mathcal{N}}^{\prime\prime},{\mathcal{N}}^{\prime})\leqslant c_{\mathcal{D}}^{\prime\prime}+c_{\mathcal{D}}^{\prime}.$
@@
Hence, there exists a tree-child SNPR sequence $\sigma$ connecting $\mathcal{N}$$^{\prime\prime}$ and $\mathcal{N}$$^{\prime}$ with $w(\sigma)\leqslant c_{\mathcal{D}}^{\prime\prime}+c_{\mathcal{D}}^{\prime}$. Moreover, since $\mathcal{N}$$^{\prime\prime}$ can be obtained from $\mathcal{N}$} by a single SNPR-, we have
$d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})\leqslant 1+c_{\mathcal{D}}^{\prime\prime}+c_{\mathcal{D}}^{\prime}=c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime},$
@@
Hence, by symmetry, there exists a tree edge $e=(u,w)$ in $\mathcal{N}$} that is not in $\mathcal{R}$}. Choose $e$ such that each directed path from $w$ to a leaf in $\mathcal{N}$} only consists of edges in $\mathcal{R}$}. Since $\mathcal{N}$} is acyclic such an edge exists. Let $D_{i}$ be the element in $\mathcal{D}$} with $i\in\{\rho,1,2,\ldots,k\}$ such that the root extension $R_{i}$ of $D_{i}$ in $\mathcal{N}$} contains $w$. By the choice of $e$, $R_{i}$ exists and each edge in $\mathcal{N}$} that lies on a directed path from $w$ to a leaf is an element of $R_{i}$. Thus, if $u=\rho$, then ${\mathcal{N}}\cong{\mathcal{N}}^{\prime}$ and the result follows as $d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})=0$. We may therefore assume that $u\neq\rho$. The next statement follows from Lemma 5(i), the additional assumption that $u\neq\rho$, and assumption (A).
13.1.
@@
By the choice of $e$, observe that the root path of $w$ consists only of $w$. In turn, $w$ has in-degree zero and out-degree zero or two in $R_{i}$. Hence, $w$ corresponds to a unique vertex, say $w_{\mathcal{D}}$, of $D_{i}$. Let $r_{w^{\prime}}$ be the vertex in $\mathcal{N}$$^{\prime}$ that $w_{\mathcal{D}}$ corresponds to. Now consider the root extension $R_{i}^{\prime}$ of $D_{i}$ in $\mathcal{N}$$^{\prime}$. Let $w^{\prime}$ be the first vertex of the root path of $r_{w^{\prime}}$. In contrast to the root path of $w$ in $R_{i}$, observe that the root path of $r_{w^{\prime}}$ may consist of more than a single vertex in which case there is a directed path of length at least one from $w^{\prime}$ to $r_{w^{\prime}}$. Since $\rho\in{\mathcal{D}}$ and $w_{\mathcal{D}}\neq\rho$, it follows that $w^{\prime}\neq\rho$. Hence, by (A), we have that $w^{\prime}$ is a leaf, or has in-degree one and out-degree two in $\mathcal{N}$$^{\prime}$. Let $v^{\prime}$ be the parent of $w^{\prime}$ in $\mathcal{N}$$^{\prime}$. As $(v^{\prime},w^{\prime})\notin{\mathcal{R}}^{\prime}$, the next statement follows from Lemma 5(i).
13.2.
@@
Let $D_{j}$ be the element in $\mathcal{D}$} with $j\in\{\rho,1,2,\ldots,k\}$ such that the root extension $R_{j}^{\prime}$ of $D_{j}$ in $\mathcal{N}$$^{\prime}$ contains $v^{\prime}$. By Lemma 5(ii), $R_{j}^{\prime}$ exists. We may have $i=j$. We next construct a digraph $\mathcal{D}$$^{\prime}$ from $\mathcal{D}$} and a network $\mathcal{N}$$^{\prime\prime}$ from $\mathcal{N}$}. After detailing the construction, we show that $\mathcal{N}$$^{\prime\prime}$ is a tree-child network that can be obtained from $\mathcal{N}$} by a single SNPR±, and that $\mathcal{D}$$^{\prime}$ is an agreement tree-child digraph for $\mathcal{N}$$^{\prime\prime}$ and $\mathcal{N}$$^{\prime}$. Guided by the second part of (13.2) and noting that, if $v^{\prime}=\rho$, then $\rho$ is a singleton component in $\mathcal{D}$}, there are three cases to consider, which are illustrated in Figure 8:
[…]
Figure 8: The three Cases (C1)–(C3) as described in the proof of Lemma 13. Blue in $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ indicates edges and vertices in $R_{i}$ and $R_{i}^{\prime}$, and green in $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ indicates edges and vertices in $R_{j}$ and $R_{j}^{\prime}$. Vertex labels $y_{1}$ and $y_{2}$ are only used to clarify the figure.
@@
$({\mathcal{R}}^{\prime}-\{R_{i}^{\prime},R_{j}^{\prime}\})\cup\{R_{ij}^{\prime}\}$
@@
We next turn to $\mathcal{D}$$^{\prime}$ and show that $\mathcal{D}$$^{\prime}$ is tree-child. Since $\mathcal{D}$} is tree-child, it follows from the definition that $\mathcal{R}$$^{\prime}$ is tree-child. Moreover, since $w^{\prime}$ has in-degree zero in $R_{i}^{\prime}$, it follows that $w^{\prime}$ has in-degree one in $R_{ij}^{\prime}$. It is now straightforward to check that $({\mathcal{R}}^{\prime}-\{R_{i}^{\prime},R_{j}^{\prime}\})\cup\{R_{ij}^{\prime}\}$ is tree-child. Hence, again by definition, $\mathcal{D}$$^{\prime}$ is also tree-child.
@@
It now follows from the construction of $\mathcal{N}$$^{\prime\prime}$ from $\mathcal{N}$} and (13.4) that $\mathcal{N}$$^{\prime\prime}$ is a phylogenetic network on $X$. For the remainder of the proof, let $D_{u}$ be the element in $\mathcal{D}$} with $u\in\{\rho,1,2,\ldots,k\}$ such that the root extension $R_{u}$ of $D_{u}$ in $\mathcal{N}$} contains $u$. By Lemma 5(ii), $R_{u}$ exists.
13.5.
@@
Again using the same notation as in the construction of $\mathcal{N}$$^{\prime\prime}$ from $\mathcal{N}$}, it follows from (13.1) that the newly added edge $(p,w)$ in $\mathcal{N}$$^{\prime\prime}$ is a tree edge. Noting that $u$ is a tree vertex by (13.1) in $\mathcal{N}$}, let $p_{u}$ be the parent of $u$, and let $c_{u}$ be the child of $u$ that is not $w$ in $\mathcal{N}$}. Observe that $(p_{u},c_{u})$ is an edge in $\mathcal{N}$$^{\prime\prime}$. Now assume that $\mathcal{N}$$^{\prime\prime}$ is not tree-child.